OpenStudy (anonymous):

What property is shown? –12 + (27 + 12) = –12 + (12 + 27)

OpenStudy (jlg030597):

Additive Property

OpenStudy (jlg030597):

take that back. Communitive

OpenStudy (anonymous):

Mhm.

OpenStudy (jlg030597):

you want proof

OpenStudy (anonymous):

No,I thought the answer was commulitive i just needed someone to confirm it.

OpenStudy (jlg030597):

is this you test or what ever 1.) What property is shown? –12 + (27 + 12) = –12 + (12 + 27) A.) associative property B.) commutative property C.) distributive property D.) opposite of a sum property 2.) Which is an example of the commutative property? A.) –13 + 27 = 27 + (–13) B.) –21 + (11 + 13) = (–21 + 11) + 13 C.) 4(–3 + 5) = 4(–3) + 4 ∙ 5 D.) –(–8 + 3) = 8 + (–3) 3.) What is the value of the expression? 19.2 + (–7 + 3.8) A.) 16 B.) 24 C.) 29 D.) 30 4.) What value of m makes the statement true? –(m + 15) = –18 + (–15) A.) –33 B.) –18 C.) 3 D.) 18 5.)Use an addition property to solve for c. 14 + (–12) = –12 + c A.) –26 B.) –14 C.) 14. D.) 26

OpenStudy (anonymous):

erm nope ,actually the next question is What is the value of the expression? –15.6 + (–3.4 + 8)

OpenStudy (jlg030597):

what answer choices.

OpenStudy (anonymous):

A. –27 B. –26 C. –11 D. –10

OpenStudy (jlg030597):

C.

OpenStudy (jlg030597):

next. Lol

OpenStudy (anonymous):

What value of q makes the statement true? –(q + 14) = –3 + (–14)

OpenStudy (anonymous):

A. –17 B. –3 C. 3 D. 17

OpenStudy (jlg030597):

working on it.

OpenStudy (jlg030597):

C

OpenStudy (anonymous):

Use an addition property to solve for a. 10 + (–15) = –15 + a A. –25 B. –10 C. –5 D. 10

OpenStudy (jlg030597):

D

OpenStudy (anonymous):

Mhm mhm.

OpenStudy (anonymous):

You got all of them right,congrats

OpenStudy (jlg030597):

Lol I love math. and thanks for the testimonial.